OpenStudy (cherryfrost27):

simone's age is 4/5 of her sisters age , what fraction of the total age of both girls is her sisters age?

alones (alones):

Um.. is that a word proble..?

OpenStudy (cherryfrost27):

yes

alones (alones):

You have to solve it out, or does it has multiple choices?

OpenStudy (cherryfrost27):

its not multiple choice..

OpenStudy (cherryfrost27):

how should i solve this? can you help me?

alones (alones):

I can try ;3

OpenStudy (cherryfrost27):

omg thank you

OpenStudy (cherryfrost27):

so what should i do first?

alones (alones):

Okay so we know that "x" is unknown..right? Let's put this together \[\frac{ 4 }{ 9 }x+x=\frac{ 9 }{ 5}x\]

alones (alones):

So I belive the asnwer should be something like \[\frac{ 4 }{ 9 }\] Since \( \frac{ 4 }{ 5 }\) over \(\ \frac{ 9 }{ 5 }\) So multiply them both..whihc would equal to \(\frac{ 4 }{ 9 }\) @cherryfrost27

alones (alones):

o.m.g pardon my spelling xD my keyboard broke.. so i have to press hard for it to work..

OpenStudy (cherryfrost27):

its okay :)